what pattern do you discover when solving even-numbered exponents with negative bases?
is it:
A.Solution is always positive.
B.Solution is always negative.
C.Solution depends on the sign of the exponent.
D.None of the Above

Answer by edjones(8007)

(-1)^6=1
(-2)^2=4
(-3)^4=81
.
(-1)^-6=1
(-2)^-4=1/81
(-3)^-2=1/9
.
A
.
Ed